# A lower bound for the Weisfeiler-Leman dimension of circulant graphs

### v1

@@ 5.8. Circulant S-rings
An S-ring ${\cal A}$ over a cyclic group $G$ is said to be circulant. Any basic set of ${\cal A}$, containing a generator of $G$, is said to be highest. The class of circulant S-rings is closed with respect to tensor product (if the underlying groups of the factors are of coprime orders) and wreath product.
As was shown in [23], any two algebraically isomorphic circulant S-rings are isomorphic, and are equal if they have the same underlying groups. Therefore, in what follows, we focus on algebraic automorphisms and isomorphisms of a circulant S-ring to itself, rather than general algebraic and combinatorial isomorphisms. The fact that any two nonisomorphic subgroups of a cyclic group $G$ have different orders, implies that if ${\cal A}$ is an S-ring over $G$ and $H$ is an ${\cal A}$-subgroup, then $H=H^{\varphi}=H^{f}$ for all algebraic automorphisms $\varphi$ of ${\cal A}$ and all normalized isomorphisms $f$ of ${\cal A}$ to itself.
@@ 5.9. Coset closure
An $S$-ring ${\cal A}$ over a group $G$ is called a coset S-ring if every its basic set is a coset by some ${\cal A}$-subgroup of $G$. The class of coset S-rings contains all group rings and is closed with respect to tensor and wreath products.
Let ${\cal A}$ be a quasidense circulant S-ring, i.e., ${\cal A}$ satisfies the following condition: if ${\mathfrak{s}}$ is an ${\cal A}$-section and the S-ring ${\cal A}_{\mathfrak{s}}$ is trivial, then the group ${\mathfrak{s}}$ is of prime order. In accordance with paper [13], the coset closure ${\cal A}_{0}$ of ${\cal A}$ is defined to be the intersection of all coset S-rings over $G$, that contain ${\cal A}$. The S-ring ${\cal A}_{0}$ is a coset one and ${\mathfrak{S}}({\cal A}_{0})={\mathfrak{S}}({\cal A})$, see Theorems 8.5 and 10.7 ibid. Denote by ${\mathfrak{S}}_{0}({\cal A})$ the set of all ${\cal A}$-sections ${\mathfrak{s}}$ such that ${\mathfrak{s}}$ is a subsection of some ${\cal A}$-section projectively equivalent to a principal ${\cal A}$-section. Then
(14) ${\mathfrak{S}}_{0}({\cal A})\subseteq\{{\mathfrak{s}}\in{\mathfrak{S}}({\cal A}):\ ({\cal A}_{0})_{\mathfrak{s}}={\mathbb{Z}}{\mathfrak{s}}\}={\mathfrak{S}}_{0}({\cal A}_{0}),$
@@ 5.10. Normal and totally normal S-rings
An S-ring ${\cal A}$ over a group $G$ is said to be normal if the subgroup $G_{right}$ is normal in $\operatorname{Aut}({\cal A})$. The class of normal S-rings contains all group rings and is closed with respect to extensions and tensor products. A complete classification of normal circulant S-rings was obtained in [9]. In particular, it was proved there that every normal circulant S-ring is cyclotomic and the following lemma holds true.
@@
Let ${\cal A}$ be a normal circulant S-ring over a group $G$ of odd order. Then
@@
    $\operatorname{WL}_{G}({\underline{X}})={\cal A}$ for any highest basic set $X\in{\cal S}({\cal A})$,
@@
    ${\cal A}=\operatorname{cyc}(M,G)$ for a unique $M\leq\operatorname{Aut}(G)$, and also $\operatorname{Aut}({\cal A})=G\rtimes M$.
@@
The assumption that the group $G$ is of odd order implies that, in terms of the paper [9], the S-ring ${\cal A}$ has trivial radical. Thus the first two statements follow from Lemma 6.7 of that paper and the proof of it, whereas the third statement is a consequence of Theorem 6.1 and Corollary 6.3 ibid. ∎
An ${\cal A}$-section ${\mathfrak{s}}$ of a circulant S-ring ${\cal A}$ is said to be normal if the S-ring ${\cal A}_{\mathfrak{s}}$ is normal. Any ${\cal A}$-section projectively equivalent to a normal section is normal. However, in general, a subsection of a normal section is not necessarily normal.
A circulant S-ring ${\cal A}$ is said to be totally normal if any subsection of a principal ${\cal A}$-section is normal, or, equivalently, any section belonging to ${\mathfrak{S}}_{0}({\cal A})$ is normal. Obviously, any circulant coset S-ring is totally normal, because the restriction of it to any principal section is a group ring (which is always normal). In general, an ${\cal A}$-section ${\mathfrak{s}}$ of a circulant S-ring ${\cal A}$ is said to be totally normal if the S-ring ${\cal A}_{\mathfrak{s}}$ is totally normal.
Any totally normal circulant S-ring ${\cal A}$ is quasidense. Indeed, otherwise, there is an ${\cal A}$-section ${\mathfrak{s}}$ of composite order and such that the S-ring ${\cal A}_{\mathfrak{s}}$ is trivial. But then ${\mathfrak{s}}$ is a subsection of a principal section. Since ${\mathfrak{s}}$ is not normal, we arrive at a contradiction.
@@ 5.11. Cayley schemes and S-rings
Let $G$ be a group. A coherent configuration ${\cal X}=(\Omega,S)$ is called a Cayley scheme over $G$ if
$\Omega=G\quad\text{and}\quad G_{right}\leq\operatorname{Aut}({\cal X}).$
Each relation $s\in S$ is the arc set of the Cayley graph $\operatorname{Cay}(G,X)$, where $X=X(s)$ is the neighborhood of the identity of $G$ in the relation $s$. The subsets $X(s)$, $s\in S$, form a partition ${\cal S}$ of the group $G$. Moreover, this partition satisfies conditions (S1), (S2), and (S3). Thus, every Cayley scheme over $G$ is associated with an S-ring ${\cal A}={\cal A}({\cal X})$ over $G$, such that ${\cal S}({\cal A})={\cal S}$.
Conversely, let ${\cal A}$ be an S-ring over $G$. For each $X\in{\cal S}({\cal A})$, denote by $s_{X}$ the arc set of the Cayley graph $\operatorname{Cay}(G,X)$. Then the relations $s_{X}$, $X\in{\cal S}({\cal A})$, form a partition $S$ of the Cartesian square $\Omega^{2}$, where $\Omega=G$. Moreover, this partition satisfies conditions (CC1), (CC2), and (CC3). Thus, every S-ring ${\cal A}$ over $G$ is associated with a Cayley scheme ${\cal X}={\cal X}({\cal A})$ over $G$, such that $S({\cal X})=S$.
@@
(15) ${\cal X}\mapsto{\cal A}({\cal X})\quad\text{and}\quad{\cal A}\mapsto{\cal X}({\cal A})$
define a one-to-one correspondence between the Cayley schemes over $G$ and the S-rings over $G$. This correspondence preserves the corresponding partial orders on the schemes and S-rings. Moreover, it is functorial in the sense that for any Cayley schemes ${\cal X}$ and ${\cal X}^{\prime}$, we have
$\operatorname{Iso}({\cal X},{\cal X}^{\prime})=\operatorname{Iso}({\cal A},{\cal A}^{\prime}),$
where ${\cal A}={\cal A}({\cal X})$ and ${\cal A}^{\prime}={\cal A}({\cal X}^{\prime})$. Furthermore, there is a one-to-one correspondence between algebraic isomorphisms, in which an algebraic isomorphism $\varphi:{\cal X}\to{\cal X}^{\prime}$, $s\mapsto s^{\prime}$, corresponds to the algebraic isomorphism $\tilde{\varphi}:{\cal A}\to{\cal A}^{\prime}$, $X(s)\mapsto X(\varphi(s))$. In what follows, we denote $\tilde{\varphi}$ just by $\varphi$.
@@
${\cal A}({\cal X}_{1}\otimes{\cal X}_{2})={\cal A}({\cal X}_{1})\otimes{\cal A}({\cal X}_{2})\quad\text{and}\quad{\cal A}({\cal X}_{1}\wr{\cal X}_{2})={\cal A}({\cal X}_{1})\wr{\cal A}({\cal X}_{2}).$
@@ 6. Multipliers of totally normal circulant S-rings
Throughout this section, ${\cal A}$ is a totally normal circulant S-ring over a (cyclic) group $G$, ${\cal S}={\cal S}({\cal A})$ and ${\mathfrak{S}}_{0}={\mathfrak{S}}_{0}({\cal A})$. We denote by ${\cal A}_{0}$ the coset closure of ${\cal A}$ and put ${\cal S}_{0}={\cal S}({\cal A}_{0})$. Finally, we denote by $\operatorname{Iso}_{\cal A}({\cal A}_{0})$ the set of all isomorphisms $f\in\operatorname{Iso}({\cal A}_{0})$ such that ${\cal A}^{f}={\cal A}$, and put
$\operatorname{Aut}_{\cal A}({\mathfrak{s}})=\operatorname{Aut}({\cal A}_{\mathfrak{s}})\cap\operatorname{Aut}({\mathfrak{s}})$
@@ 6.1. Outer multipliers
Suppose that for every section ${\mathfrak{s}}\in{\mathfrak{S}}_{0}$, we are given a coset ${\mathfrak{c}}_{\mathfrak{s}}=\sigma_{\mathfrak{s}}\cdot\operatorname{Aut}_{\cal A}({\mathfrak{s}})=\operatorname{Aut}_{\cal A}({\mathfrak{s}})\cdot\sigma_{\mathfrak{s}}$, where $\sigma_{\mathfrak{s}}\in\operatorname{Aut}({\mathfrak{s}})$. Following [14], the family
@@
is called an outer ${\cal A}$-multiplier if the following conditions are satisfied for all sections ${\mathfrak{s}},{\mathfrak{t}}\in{\mathfrak{S}}_{0}$:
@@
The set of all outer ${\cal A}$-multipliers forms a group $\operatorname{{\mathfrak{C}}}({\cal A})$ with respect to componentwise multiplication. By [14, Theorem 3.4], there is an isomorphism ${\mathfrak{c}}\mapsto\varphi_{\mathfrak{c}}$, ${\mathfrak{c}}\in\operatorname{{\mathfrak{C}}}({\cal A})$, from the group $\operatorname{{\mathfrak{C}}}({\cal A})$ onto the group of all algebraic automorphisms of the S-ring ${\cal A}$. Here the algebraic automorphism $\varphi=\varphi_{\mathfrak{c}}$ is defined as follows. Let $X\in{\cal S}({\cal A})$ and let ${\mathfrak{s}}={\mathfrak{s}}_{X}$ the principal section associated with $X$. Then we put $X^{\varphi}$ to be the $\pi_{\mathfrak{s}}$-preimage of the set $(X_{\mathfrak{s}})^{\sigma_{\mathfrak{s}}}$, see formula (11). Thus,
@@ 6.2. Inner multipliers
is called an inner ${\cal A}$-multiplier if the conditions (17) are satisfied for the cosets ${\mathfrak{c}}_{\mathfrak{s}}$ replaced by the singletons $\{{\mathfrak{m}}_{\mathfrak{s}}\}$. The restriction ${\mathfrak{m}}_{{\mathfrak{T}}}$ of the inner multiplier ${\mathfrak{m}}$ to a subset ${\mathfrak{T}}\subseteq{\mathfrak{S}}_{0}$ is defined to be the family $\{{\mathfrak{m}}_{\mathfrak{s}}:\ {\mathfrak{s}}\in{\mathfrak{T}}\}$.
The set of all inner ${\cal A}$-multipliers forms a subgroup $\operatorname{{\mathfrak{M}}}({\cal A})$ of the direct product of the groups $\operatorname{Aut}({\mathfrak{s}})$, ${\mathfrak{s}}\in{\mathfrak{S}}_{0}$. For a subgroup ${\mathfrak{M}}\leq{\mathfrak{M}}({\cal A})$ and a set ${\mathfrak{T}}\subseteq{\mathfrak{S}}_{0}$, we put ${\mathfrak{M}}_{\mathfrak{T}}=\{{\mathfrak{m}}_{\mathfrak{T}}:\ {\mathfrak{m}}\in{\mathfrak{M}}\}$ and abbreviate ${\mathfrak{M}}_{\mathfrak{s}}:={\mathfrak{M}}_{\{{\mathfrak{s}}\}}$ for ${\mathfrak{s}}\in{\mathfrak{S}}_{0}$. Clearly, ${\mathfrak{M}}_{\mathfrak{T}}$ is a group with respect to componentwise product. For any inner ${\cal A}$-multiplier ${\mathfrak{m}}$ and any outer ${\cal A}$-multiplier ${\mathfrak{c}}$, we write ${\mathfrak{m}}_{\mathfrak{T}}\leq{\mathfrak{c}}_{\mathfrak{T}}$ if ${\mathfrak{m}}_{\mathfrak{s}}\in\ {\mathfrak{c}}_{\mathfrak{s}}$ for all ${\mathfrak{s}}\in{\mathfrak{T}}$. When ${\mathfrak{T}}={\mathfrak{S}}_{0}$, we abbreviate ${\mathfrak{m}}\leq{\mathfrak{c}}$.
@@
For each inner ${\cal A}$-multiplier ${\mathfrak{m}}\in\operatorname{{\mathfrak{M}}}({\cal A})$, there exists a normalized isomorphism $f_{\mathfrak{m}}\in\operatorname{Iso}_{\cal A}({\cal A}_{0})$ such that for all $X\in{\cal S}$, we have
@@
By [14, Lemma 5.2], the family of the singletons $\{{\mathfrak{m}}_{\mathfrak{s}}\}$, ${\mathfrak{s}}\in{\mathfrak{S}}_{0}$, is an outer ${\cal A}_{0}$-multiplier. By [14, Theorem 3.4], there exists a uniquely determined algebraic automorphism $\varphi$ of the S-ring ${\cal A}_{0}$ such that
(21) ${X_{\mathfrak{s}}}^{\varphi_{\mathfrak{s}}}={X_{\mathfrak{s}}}^{{\mathfrak{m}}_{\mathfrak{s}}},\qquad X\in{\cal S},$
where ${\mathfrak{s}}={\mathfrak{s}}_{X}$. By [13, Theorem 9.1], every algebraic automorphism of ${\cal A}_{0}$ is induced by an isomorphism. Consequently, there exists a normalized isomorphism $f\in\operatorname{Iso}({\cal A}_{0},\varphi)$. By virtue of (21), we have
@@
To prove that $f\in\operatorname{Iso}_{\cal A}({\cal A}_{0})$, we need to verify that $X^{f}\in{\cal S}$ for all $X\in{\cal S}$. The principal section ${\mathfrak{s}}={\mathfrak{s}}_{X}$ belongs to the set ${\mathfrak{S}}_{0}$. By formula (11),

### v2

@@ 5.8. Circulant S-rings
An S-ring $\mathcal{A}$} over a cyclic group $G$ is said to be circulant. Any basic set of $\mathcal{A}$}, containing a generator of $G$, is said to be highest. The class of circulant S-rings is closed with respect to tensor product (if the underlying groups of the factors are of coprime orders) and wreath product.
As was shown in [23], any two algebraically isomorphic circulant S-rings are isomorphic, and are equal if they have the same underlying groups. Therefore, in what follows, we focus on algebraic automorphisms and isomorphisms of a circulant S-ring to itself, rather than general algebraic and combinatorial isomorphisms. The fact that any two nonisomorphic subgroups of a cyclic group $G$ have different orders, implies that if $\mathcal{A}$} is an S-ring over $G$ and $H$ is an $\mathcal{A}$}-subgroup, then $H=H^{\varphi}=H^{f}$ for all algebraic automorphisms $\varphi$ of $\mathcal{A}$} and all normalized isomorphisms $f$ of $\mathcal{A}$} to itself.
@@ 5.9. Coset closure
An $S$-ring $\mathcal{A}$} over a group $G$ is called a coset S-ring if every its basic set is a coset by some $\mathcal{A}$}-subgroup of $G$. The class of coset S-rings contains all group rings and is closed with respect to tensor and wreath products.
Let $\mathcal{A}$} be a quasidense circulant S-ring, i.e., $\mathcal{A}$} satisfies the following condition: if $\mathfrak{s}$} is an $\mathcal{A}$}-section and the S-ring ${\mathcal{A}}_{\mathfrak{s}}$ is trivial, then the group $\mathfrak{s}$} is of prime order. In accordance with paper [13], the coset closure ${\mathcal{A}}_{0}$ of $\mathcal{A}$} is defined to be the intersection of all coset S-rings over $G$, that contain $\mathcal{A}$}. The S-ring ${\mathcal{A}}_{0}$ is a coset one and ${\mathfrak{S}}({\mathcal{A}}_{0})={\mathfrak{S}}({\mathcal{A}})$, see Theorems 8.5 and 10.7 ibid. Denote by ${\mathfrak{S}}_{0}({\mathcal{A}})$ the set of all $\mathcal{A}$}-sections $\mathfrak{s}$} such that $\mathfrak{s}$} is a subsection of some $\mathcal{A}$}-section projectively equivalent to a principal $\mathcal{A}$}-section. Then
(14) ${\mathfrak{S}}_{0}({\mathcal{A}})\subseteq\{{\mathfrak{s}}\in{\mathfrak{S}}({\mathcal{A}}):\ ({\mathcal{A}}_{0})_{\mathfrak{s}}={\mathbb{Z}}{\mathfrak{s}}\}={\mathfrak{S}}_{0}({\mathcal{A}}_{0}),$
@@ 5.10. Normal and totally normal S-rings
An S-ring $\mathcal{A}$} over a group $G$ is said to be normal if the subgroup $G_{right}$ is normal in $\operatorname{Aut}({\mathcal{A}})$. The class of normal S-rings contains all group rings and is closed with respect to extensions and tensor products. A complete classification of normal circulant S-rings was obtained in [9]. In particular, it was proved there that every normal circulant S-ring is cyclotomic and the following lemma holds true.
@@
Let $\mathcal{A}$} be a normal circulant S-ring over a group $G$ of odd order. Then
@@
    $\operatorname{WL}_{G}({\underline{X}})={\mathcal{A}}$ for any highest basic set $X\in{\mathcal{S}}({\mathcal{A}})$,
@@
    ${\mathcal{A}}=\operatorname{cyc}(M,G)$ for a unique $M\leq\operatorname{Aut}(G)$, and also $\operatorname{Aut}({\mathcal{A}})=G\rtimes M$.
@@
The assumption that the group $G$ is of odd order implies that, in terms of the paper [9], the S-ring $\mathcal{A}$} has trivial radical. Thus the first two statements follow from Lemma 6.7 of that paper and the proof of it, whereas the third statement is a consequence of Theorem 6.1 and Corollary 6.3 ibid. ∎
An $\mathcal{A}$}-section $\mathfrak{s}$} of a circulant S-ring $\mathcal{A}$} is said to be normal if the S-ring ${\mathcal{A}}_{\mathfrak{s}}$ is normal. Any $\mathcal{A}$}-section projectively equivalent to a normal section is normal. However, in general, a subsection of a normal section is not necessarily normal.
A circulant S-ring $\mathcal{A}$} is said to be totally normal if any subsection of a principal $\mathcal{A}$}-section is normal, or, equivalently, any section belonging to ${\mathfrak{S}}_{0}({\mathcal{A}})$ is normal. Obviously, any circulant coset S-ring is totally normal, because the restriction of it to any principal section is a group ring (which is always normal). In general, an $\mathcal{A}$}-section $\mathfrak{s}$} of a circulant S-ring $\mathcal{A}$} is said to be totally normal if the S-ring ${\mathcal{A}}_{\mathfrak{s}}$ is totally normal.
Any totally normal circulant S-ring $\mathcal{A}$} is quasidense. Indeed, otherwise, there is an $\mathcal{A}$}-section $\mathfrak{s}$} of composite order and such that the S-ring ${\mathcal{A}}_{\mathfrak{s}}$ is trivial. But then $\mathfrak{s}$} is a subsection of a principal section. Since $\mathfrak{s}$} is not normal, we arrive at a contradiction.
@@ 5.11. Cayley schemes and S-rings
Let $G$ be a group. A coherent configuration ${\mathcal{X}}=(\Omega,S)$ is called a Cayley scheme over $G$ if
$\Omega=G\quad\text{and}\quad G_{right}\leq\operatorname{Aut}({\mathcal{X}}).$
Each relation $s\in S$ is the arc set of the Cayley graph $\operatorname{Cay}(G,X)$, where $X=X(s)$ is the neighborhood of the identity of $G$ in the relation $s$. The subsets $X(s)$, $s\in S$, form a partition $\mathcal{S}$} of the group $G$. Moreover, this partition satisfies conditions (S1), (S2), and (S3). Thus, every Cayley scheme over $G$ is associated with an S-ring ${\mathcal{A}}={\mathcal{A}}({\mathcal{X}})$ over $G$, such that ${\mathcal{S}}({\mathcal{A}})={\mathcal{S}}$.
Conversely, let $\mathcal{A}$} be an S-ring over $G$. For each $X\in{\mathcal{S}}({\mathcal{A}})$, denote by $s_{X}$ the arc set of the Cayley graph $\operatorname{Cay}(G,X)$. Then the relations $s_{X}$, $X\in{\mathcal{S}}({\mathcal{A}})$, form a partition $S$ of the Cartesian square $\Omega^{2}$, where $\Omega=G$. Moreover, this partition satisfies conditions (CC1), (CC2), and (CC3). Thus, every S-ring $\mathcal{A}$} over $G$ is associated with a Cayley scheme ${\mathcal{X}}={\mathcal{X}}({\mathcal{A}})$ over $G$, such that $S({\mathcal{X}})=S$.
@@
(15) ${\mathcal{X}}\mapsto{\mathcal{A}}({\mathcal{X}})\quad\text{and}\quad{\mathcal{A}}\mapsto{\mathcal{X}}({\mathcal{A}})$
define a one-to-one correspondence between the Cayley schemes over $G$ and the S-rings over $G$. This correspondence preserves the corresponding partial orders on the schemes and S-rings. Moreover, it is functorial in the sense that for any Cayley schemes $\mathcal{X}$} and $\mathcal{X}$$^{\prime}$, we have
$\operatorname{Iso}({\mathcal{X}},{\mathcal{X}}^{\prime})=\operatorname{Iso}({\mathcal{A}},{\mathcal{A}}^{\prime}),$
where ${\mathcal{A}}={\mathcal{A}}({\mathcal{X}})$ and ${\mathcal{A}}^{\prime}={\mathcal{A}}({\mathcal{X}}^{\prime})$. Furthermore, there is a one-to-one correspondence between algebraic isomorphisms, in which an algebraic isomorphism $\varphi:{\mathcal{X}}\to{\mathcal{X}}^{\prime}$, $s\mapsto s^{\prime}$, corresponds to the algebraic isomorphism $\tilde{\varphi}:{\mathcal{A}}\to{\mathcal{A}}^{\prime}$, $X(s)\mapsto X(\varphi(s))$. In what follows, we denote $\tilde{\varphi}$ just by $\varphi$.
@@
${\mathcal{A}}({\mathcal{X}}_{1}\otimes{\mathcal{X}}_{2})={\mathcal{A}}({\mathcal{X}}_{1})\otimes{\mathcal{A}}({\mathcal{X}}_{2})\quad\text{and}\quad{\mathcal{A}}({\mathcal{X}}_{1}\wr{\mathcal{X}}_{2})={\mathcal{A}}({\mathcal{X}}_{1})\wr{\mathcal{A}}({\mathcal{X}}_{2}).$
@@ 6. Multipliers of totally normal circulant S-rings
Throughout this section, $\mathcal{A}$} is a totally normal circulant S-ring over a (cyclic) group $G$, ${\mathcal{S}}={\mathcal{S}}({\mathcal{A}})$ and ${\mathfrak{S}}_{0}={\mathfrak{S}}_{0}({\mathcal{A}})$. We denote by ${\mathcal{A}}_{0}$ the coset closure of $\mathcal{A}$} and put ${\mathcal{S}}_{0}={\mathcal{S}}({\mathcal{A}}_{0})$. Finally, we denote by $\operatorname{Iso}_{\mathcal{A}}({\mathcal{A}}_{0})$ the set of all isomorphisms $f\in\operatorname{Iso}({\mathcal{A}}_{0})$ such that ${\mathcal{A}}^{f}={\mathcal{A}}$, and put
$\operatorname{Aut}_{\mathcal{A}}({\mathfrak{s}})=\operatorname{Aut}({\mathcal{A}}_{\mathfrak{s}})\cap\operatorname{Aut}({\mathfrak{s}})$
@@ 6.1. Outer multipliers
Suppose that for every section ${\mathfrak{s}}\in{\mathfrak{S}}_{0}$, we are given a coset ${\mathfrak{c}}_{\mathfrak{s}}=\sigma_{\mathfrak{s}}\cdot\operatorname{Aut}_{\mathcal{A}}({\mathfrak{s}})=\operatorname{Aut}_{\mathcal{A}}({\mathfrak{s}})\cdot\sigma_{\mathfrak{s}}$, where $\sigma_{\mathfrak{s}}\in\operatorname{Aut}({\mathfrak{s}})$. Following [14], the family
@@
is called an outer $\mathcal{A}$}-multiplier if the following conditions are satisfied for all sections ${\mathfrak{s}},{\mathfrak{t}}\in{\mathfrak{S}}_{0}$:
@@
The set of all outer $\mathcal{A}$}-multipliers forms a group $\operatorname{{\mathfrak{C}}}({\mathcal{A}})$ with respect to componentwise multiplication. By [14, Theorem 3.4], there is an isomorphism ${\mathfrak{c}}\mapsto\varphi_{\mathfrak{c}}$, ${\mathfrak{c}}\in\operatorname{{\mathfrak{C}}}({\mathcal{A}})$, from the group $\operatorname{{\mathfrak{C}}}({\mathcal{A}})$ onto the group of all algebraic automorphisms of the S-ring $\mathcal{A}$}. Here the algebraic automorphism $\varphi=\varphi_{\mathfrak{c}}$ is defined as follows. Let $X\in{\mathcal{S}}({\mathcal{A}})$ and let ${\mathfrak{s}}={\mathfrak{s}}_{X}$ the principal section associated with $X$. Then we put $X^{\varphi}$ to be the $\pi_{\mathfrak{s}}$-preimage of the set $(X_{\mathfrak{s}})^{\sigma_{\mathfrak{s}}}$, see formula (11). Thus,
@@ 6.2. Inner multipliers
is called an inner $\mathcal{A}$}-multiplier if the conditions (17) are satisfied for the cosets ${\mathfrak{c}}_{\mathfrak{s}}$ replaced by the singletons $\{{\mathfrak{m}}_{\mathfrak{s}}\}$. The restriction ${\mathfrak{m}}_{{\mathfrak{T}}}$ of the inner multiplier $\mathfrak{m}$} to a subset ${\mathfrak{T}}\subseteq{\mathfrak{S}}_{0}$ is defined to be the family $\{{\mathfrak{m}}_{\mathfrak{s}}:\ {\mathfrak{s}}\in{\mathfrak{T}}\}$.
The set of all inner $\mathcal{A}$}-multipliers forms a subgroup $\operatorname{{\mathfrak{M}}}({\mathcal{A}})$ of the direct product of the groups $\operatorname{Aut}({\mathfrak{s}})$, ${\mathfrak{s}}\in{\mathfrak{S}}_{0}$. For a subgroup ${\mathfrak{M}}\leq{\mathfrak{M}}({\mathcal{A}})$ and a set ${\mathfrak{T}}\subseteq{\mathfrak{S}}_{0}$, we put ${\mathfrak{M}}_{\mathfrak{T}}=\{{\mathfrak{m}}_{\mathfrak{T}}:\ {\mathfrak{m}}\in{\mathfrak{M}}\}$ and abbreviate ${\mathfrak{M}}_{\mathfrak{s}}:={\mathfrak{M}}_{\{{\mathfrak{s}}\}}$ for ${\mathfrak{s}}\in{\mathfrak{S}}_{0}$. Clearly, ${\mathfrak{M}}_{\mathfrak{T}}$ is a group with respect to componentwise product. For any inner $\mathcal{A}$}-multiplier $\mathfrak{m}$} and any outer $\mathcal{A}$}-multiplier $\mathfrak{c}$}, we write ${\mathfrak{m}}_{\mathfrak{T}}\leq{\mathfrak{c}}_{\mathfrak{T}}$ if ${\mathfrak{m}}_{\mathfrak{s}}\in\ {\mathfrak{c}}_{\mathfrak{s}}$ for all ${\mathfrak{s}}\in{\mathfrak{T}}$. When ${\mathfrak{T}}={\mathfrak{S}}_{0}$, we abbreviate ${\mathfrak{m}}\leq{\mathfrak{c}}$.
@@
For each inner $\mathcal{A}$}-multiplier ${\mathfrak{m}}\in\operatorname{{\mathfrak{M}}}({\mathcal{A}})$, there exists a normalized isomorphism $f_{\mathfrak{m}}\in\operatorname{Iso}_{\mathcal{A}}({\mathcal{A}}_{0})$ such that for all $X\in{\mathcal{S}}$, we have
@@
By [14, Lemma 5.2], the family of the singletons $\{{\mathfrak{m}}_{\mathfrak{s}}\}$, ${\mathfrak{s}}\in{\mathfrak{S}}_{0}$, is an outer ${\mathcal{A}}_{0}$-multiplier. By [14, Theorem 3.4], there exists a uniquely determined algebraic automorphism $\varphi$ of the S-ring ${\mathcal{A}}_{0}$ such that
(21) ${X_{\mathfrak{s}}}^{\varphi_{\mathfrak{s}}}={X_{\mathfrak{s}}}^{{\mathfrak{m}}_{\mathfrak{s}}},\qquad X\in{\mathcal{S}},$
where ${\mathfrak{s}}={\mathfrak{s}}_{X}$. By [13, Theorem 9.1], every algebraic automorphism of ${\mathcal{A}}_{0}$ is induced by an isomorphism. Consequently, there exists a normalized isomorphism $f\in\operatorname{Iso}({\mathcal{A}}_{0},\varphi)$. By virtue of (21), we have
@@
To prove that $f\in\operatorname{Iso}_{\mathcal{A}}({\mathcal{A}}_{0})$, we need to verify that $X^{f}\in{\mathcal{S}}$ for all $X\in{\mathcal{S}}$. The principal section ${\mathfrak{s}}={\mathfrak{s}}_{X}$ belongs to the set ${\mathfrak{S}}_{0}$. By formula (11),